OpenStudy (anonymous):

What is the length of the hypotenuse of a right triangle with legs of lengths 7cm and 24cm? show work please

OpenStudy (anonymous):

use pythagores (spell) theorem: \[hypotenuse = \sqrt{(7)^2+(24)^2} = 25 \space cm\]

OpenStudy (anonymous):

25

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

hypothenuse h=sqrt(x^2 +y^2)

OpenStudy (anonymous):

yw good luck